OpenStudy (anonymous):

HELP.. differential problem y' = (3x^2-1)/(3+2y)

OpenStudy (anonymous):

no, i need the value of y

OpenStudy (anonymous):

difeq right? multiply by \[3+2y\] get \[(2y+3 )dy = x^3+1\] integrate both sides get \[y^2+3y=x^4+x+c\]

OpenStudy (anonymous):

that was wrong

OpenStudy (anonymous):

\[(2y+3 )dy = 3x^2+1\] \[y^2+3x=x^3+x+c\] is better, sorry about that

OpenStudy (anonymous):

well except the plus x should be minus x i need more coffee, but i hope you get the idea. then solve for y

OpenStudy (anonymous):

thanks, this is really helpfull

OpenStudy (anonymous):

solving is ok, right? use quadratic formula